Sample Mean Range 16 8 12 13 6 6 4 15 10 10 11 10 16 13 14 0 14 17 11 12 IU 12 12 11 13 14 12 7 8 13 9 11 14 7 13 15 10 15 SuQuestion pertains to the x-chart part
Answer #1

Answer a:

R= Max- Min X-bar= average of sample
8 16
13 12
6 6
10 15
11 10
16 10
14 13
17 14
12 11
12 12
14 13
8 7
11 9
13 7
15 10
12.000 11.000
R-bar (Average of above values) X-bar-bar (average of above values)

D4, D3 and A2 are taken from table of factors computing 3 sigma control limits, sample size=5

Range

D4 (n=5)= 2.115
D3 (n=5)= 0
control limits for Range,
CL or R-bar= 12.00
UCL=R-bar*D4 25.38
LCL=R-bar*D3 0.00

Xbar

A2 (n=5) 0.577
control limits for X bar,
CL or Xbarbar 11.00
UCL=Xbarbar+ (A2)*Rbar 17.92
LCL=Xbarbar- (A2)*Rbar 4.08